OpenStudy (anonymous):

HELP ASAP !!!!!! Suppose that there are two types of tickets to a show: advance and same-day. The combined cost of one advance ticket and one same-day ticket is $35. For one performance, 20 advance tickets and 35 same-day tickets were sold. The total amount paid for the tickets was $925. What was the price of each kind of ticket?

OpenStudy (anonymous):

@iLorita

OpenStudy (anonymous):

https://answers.yahoo.com/question/index?qid=20130927135456AAxe3kh Here is a similar question...just follow the explanations yet change the numbers...i hope this helps.

OpenStudy (anonymous):

were did the 5x+6y=215 come from

OpenStudy (anonymous):

SOMEBODY HELP ME !!!!!!!!!!!!!!!!!!! @inowalst @inkyvoyd @logan420

OpenStudy (anonymous):

@pinklion23

OpenStudy (pinklion23):

im here

OpenStudy (pinklion23):

Set up a set of 2 equations in 2 variables - Let x = advance ticket cost and y = same-day tickets cost Then x+y=40 and 25x+30y=1075 >>>>> 5x+6y=215 x=40-y, so 5(40-y)+6y=215 200-5y+6y=215 y=15 >>>> x=25 5(25) + 6(15) = 125+90 = 215

OpenStudy (pinklion23):

hope it helped :)

OpenStudy (anonymous):

i don't understand where 5x+6y=215 came from

OpenStudy (pinklion23):

this confuses me too but this is what this guy said so :/